# Review: Quantum Metrology and Sensing with Many-Body Systems

### v2

@@ VII Resource analysis for critical quantum sensing
As discussed above, various forms of criticality can achieve quantum enhanced sensitivity. The resource for achieving such enhanced precision is the system size $N$ and the figure of merit is the QFI. However, in some cases the time $T$ needed for accomplishing the sensing task also has to be considered as a resource. This might be due to practical issues, such as limited coherence time, or size dependent $T$ which makes the sensing time extremely long for large system sizes. To account for such complexity, one can use normalized QFI as a figure of merit, namely $I_{Q}/T$ Albarelli et al. (2018); Rossi et al. (2020); Chin et al. (2012); Chaves et al. (2013); Brask et al. (2015); Smirne et al. (2016). For instance, in the case of many-body ground state sensing, one may consider an adiabatic approach for generating the complex ground state. This can be achieved by slowly evolving the system initialized in the ground state of a simpler Hamiltonian which is transformed into the desired Hamiltonian over the time interval $T$. If the transformation is slow enough, namely $T$ is large, the quantum state remains in the ground state of the instant Hamiltonian throughout the evolution. This guarantees that at the end of evolution when the Hamiltonian takes the final form the system is in the desired ground state. When the system is adiabatically driven towards a second-order quantum phase transition, the required time $T$ scales as $T{\sim}N^{z}$ Rams et al. (2018), where $z$ is the dynamical critical exponent showing how the energy gap $\Delta E$ vanishes at the critical point by increasing the system size $\Delta E_{\rm min}\sim 1/N^{z}$. Incorporating this into the scaling of the QFI across second-order quantum phase transitions in Eq. (39) leads to
@@
This clearly shows that the scaling diminishes as time is included in the resource analysis.
@@
Similar results can be found for different quantum sensing strategies where the required time depends on the system size. For instance, in Ref. Montenegro et al. (2023) quantum enhanced sensitivity, i.e. $\beta{>}1$, is achieved for the steady state at the boundary time crystal transition when the figure of merit is the QFI $I_{Q}$. However, if one changes their figure of merit to normalized QFI, i.e. $I_{Q}/T$, with $T$ being the time that the system reaches its steady state, the scaling with respect to system size is reduced to standard limit.
@@
The time needed for accomplishing the sensing task, is directly related to the strategy that one exploits. In Ref. Gietka et al. (2021), the authors show that by adiabatic evolution, one cannot get advantage for quantum sensing. However, the proof relies on a specific choice of the adiabatic evolution. In their formalism the Hamiltonian is written as $H{=}\theta H_{0}{+}H_{1}(t)$, where $H_{0}$ and $H_{1}$ do not depend on the target parameter $\theta$. In this scenario, $H_{0}$ is always on and the time dependent Hamiltonian $H(t)$ varies slowly. The resulted dynamics encodes the information about $\theta$ into the state of the system whose initial state does not depend on $\theta$. In this case, as proved in Ref. Gietka et al. (2021), there is no quantum advantage. Nonetheless, one may consider to prepare the probe in the ground state of $H_{1}$ and as $H_{1}$ is turned off, the probe is put under the action of a general $H_{0}(\theta)$ adiabatically to prepare the probe in the ground state of a general $H_{0}(\theta)$. In this situation, obtaining quantum advantage is possible. For instance, in the case of Stark systems, presented in section V.2, it has been shown that regardless of the number of particles and the range of the interaction, one always has $\beta-z{>}1$ ensuring that the quantum enhanced sensitivity is always achievable for the Stark probes He et al. (2023); Yousefjani et al. (2023a).
@@ VIII Non-Equilibrium Quantum Sensing: Dynamical Quantum Phase Transitions
The footprints of the ground state quantum criticality in equilibrium can also be observed in non-equilibrium dynamics. For a closed system described by a time-independent Hamiltonian $H$, its time evolution which is described by Schrödinger equation as
$|\Psi(t)\rangle=e^{-iHt}|\Psi(0)\rangle,$ (63)
where $|\Psi(0)\rangle$ is the initial state. There are two different notions of dynamical quantum phase transitions which we will discuss separately in the following subsections.
@@ VIII.1 Dynamical quantum phase transitions quantified by Loschmidt echo
By considering the projection of the time-evolved state with the initial state, one can defined the Loschmidt echo, ${\cal L}(t)=|\langle\Psi(0)|\Psi(t)\rangle|^{2}$, interpreted as the return probability. One can define the rate of return probability ${\cal G}$, as
${\cal G}=-\lim_{N\to\infty}\frac{1}{N}\log{\cal}L(t).$ (64)
It is discovered that if $|\Psi(0)\rangle$ and $|\Psi(t)\rangle$ belongs to two-different equilibrium phases, then the ${\cal G}(t)$ can be non-analytic at certain time $t{=}t_{c}$. On the other hands, if $|\Psi(0)\rangle$ and $|\Psi(t)\rangle$ belongs to the same equilibrium phase then ${\cal G}(t)$ is a regular function of time Heyl (2018). There are a few peculiarities to these cases Jafari and Johannesson (2017) as well as extension in non-Hermitian systems Zhou et al. (2018); Jing et al. (2024). Such dynamical phase transitions have been experimentally realized in large quantum simulators (Zhang et al., 2017) and has also been proposed to provide quantum enhancement in parameter estimation González et al. (2022), where the authors have proposed a negatively charged Nitrogen-Vacancy (NV-) center in a nearby Carbon-13 nuclear spins as a platform for studying dynamical phase transitions. First, they have shown that the system undergoes a dynamical phase transition by a suitable choice of Gaussian time-varying magnetic field. Then it is shown that the Fisher information, for estimating coupling strength between the two Carbon-13 nuclear spins, improves at $t{=}t_{c}$ as compared to the regular times.
@@ VIII.2 Dynamical quantum phase transitions quantified by a time-averaged order parameter
A second category of dynamical phase transitions in many-body systems is detected by time-averaged order parameter ${\cal O}(t)$. The time average of the order parameter in the evolved state $|\Psi(t)\rangle$ is given by
@@
The order parameter distinguishes an ordered phase when $\overline{\langle{\cal O}(t)\rangle}\neq 0$ from a disordered phase when $\overline{\langle{\cal O}(t)\rangle}=0$. Such phase transitions have been identified for quantum-enhanced sensing Guan and Lewis-Swan (2021). In this paper, the authors have studied a fully connected spin model under sudden quenching. They focused on the LMG model which is given by
$H_{LMG}{=}-\frac{\chi}{N}S^{2}_{z}-\Omega S_{x}-\omega S_{z}$ (66)
where $S_{\alpha}{=}\sum_{i=1}^{N}\sigma_{\alpha,i}/2$ are defined as collective spin operators and $\sigma_{\alpha,i}$ are the Pauli matrices for the $i$th spin-1/2 particle. The ground state criticality of the model occurs at $\Omega_{c}/\chi{=}0.5$ and $\omega_{c}/\chi{=}0$. The Fisher information in the long time state behaves as an order parameter and shows peak at the critical point. Moreover, the scaling of the long-time QFI at the critical point as a function of system size $N$ reaches to a sub-Heisenberg limit for the transverse field, $I^{t\to\infty}_{Q,B_{x}}\sim N^{1.5}$ and $I^{t\to\infty}_{Q,B_{z}}\sim N^{1.75}$ for the longitudinal field.
@@ IX Non-Equilibrium Quantum Sensing: Discrete Time Crystal Phase Transitions
Apart from dynamical phase transitions, discussed in Sec. VIII, there are other categories of phase transitions which occur through non-equilibrium dynamics of many-body systems. An important class of such systems are discrete-time crystals (DTCs) which are defined in an isolated non-equilibrium quantum many-body system that undergoes periodic drives. Breaking the time transnational symmetry results in the emergence of a new phase of matter known as the DTC phase with indefinite persistent oscillations Sacha (2015); Khemani et al. (2016); Else et al. (2016); Yoshinaga et al. (2022). Recently the long-range spatial- and time-ordered dynamics in the DTC phase have been harnessed in measuring an AC filed Iemini et al. (2024b); Gribben et al. (2024). The optimal performance for long periods along with intrinsic robustness to the imperfections in the sensing protocol based on DTC phase allows one to hit the quantum standard limit in estimating the AC field. However, obtaining a true sensing enhancement based on a resource such as system size requires a wised mechanism for establishing DTC. In Ref. Yousefjani et al. (2024b) the proposed mechanism results in a stable DTC with period-doubling oscillations that persist indefinitely even in finite-size systems. The time-dependent Hamiltonian that governs the dynamic in the considered spin-$1/2$ chain is
$H(t)=JH_{I}+\sum_{n}\delta(t-nT)H_{P},$ (67)
$H_{I}=\sum_{j=1}^{L-1}j\sigma^{z}_{j}\sigma^{z}_{j+1},\quad H_{P}=\Phi\sum_{j=1}^{L}\sigma^{x}_{j}.$ (68)
@@
Here $J$ is the spin exchange coupling, and $\sigma^{x,y,z}_{j}$ are the Pauli operators. For one period of evolution, the Floquet unitary operator reads
$U_{F}(\Omega,\varepsilon)=e^{-iH_{P}}e^{-i\Omega H_{I}},$ (69)
@@
here $\Omega{=}JT$, and $\Phi$ is tuned to be $\Phi{=}(1{-}\varepsilon)\frac{\pi}{2}$, with $\varepsilon$ as deviation from a $\pi/2$ $x$-rotation. While setting $\Omega{=}\pi/2$ results in a stable period doubling DTC that is robust against arbitrary imperfection $\varepsilon$ in the rotating pulse, the system goes through a sharp second-order phase transition as the spin exchange coupling, namely $\Omega$ varies from $\frac{\pi}{2}$, denoting this variation as $\omega{=}|\Omega-\frac{\pi}{2}|$. Relying on this transition, a DTC quantum sensor has been devised to sense $\omega$ with benefits from multiple features. As it is clear from Fig. 11(a), in the DTC phase the QFI provides a plateau whose value depends on period cycles $n$, and in the non-DTC region it shows nontrivial and fast oscillations. By approaching the transition point, denoted by $\omega_{\max}$ (dashed line), the QFI shows a clear peak at all stroboscopic times. The extreme sensitivity to the exchange coupling across the whole DTC phase ($\omega{<}\omega_{\max}$) as well as at the transition point ($\omega_{\max}$) providing quantum-enhanced sensitivity as $I_{Q}{\propto}L^{\beta}$ with $\beta{>}3$ (see Fig.11 (b)). The obtained quantum enhancement is independent of the initial state and can be boosted by increasing imperfection in the pulse to a certain value.
@@
Figure 11: DTC quantum sensor; (a) The QFI versus $\omega$ in stroboscopic times in a system of size $L{=}30$. The onset of the phase transition is determined by $\omega{=}\omega_{\max}$ (dashed line), the point where QFI peaks in different period cycles $n$’s. (b) The values of the QFI after $n=2$ in DTC phase (for $\omega{=}10^{-4}\pi/2$) and at transition points ($\omega{=}\omega_{\max}$) versus $L$. The numerical simulation (NS) is well-mapped by a function as $I_{Q}{\propto}L^{\beta}$ (solid lines) with $\beta{>}3$. Results are obtained for $\varepsilon{=}0.01$. Figure is adapted from Yousefjani et al. (2024b).

### v3

@@ VII Resource Analysis for Critical Quantum Sensing
As discussed above, various forms of criticality can achieve quantum-enhanced sensitivity. The resource for achieving such enhanced precision is the system size $N$ and the figure of merit is the QFI. However, in some cases the time $t$ needed for accomplishing the sensing task also has to be considered as a resource. This might be due to practical issues, such as limited coherence time, or size dependent $t$ which can make the sensing time extremely long for large system sizes. To account for such complexity, one can use normalized QFI as a figure of merit, namely $I_{Q}/t$ Albarelli et al. (2018); Rossi et al. (2020); Chin et al. (2012); Chaves et al. (2013); Brask et al. (2015); Smirne et al. (2016). This can be understood by considering the total time $t_{\mathrm{tot}}$ that is needed for collecting the data through probe preparation and measurement with $M$ repetitions. The number of measurements that are possible within the available total time is $M{=}t_{\mathrm{tot}}/t$. Now from Eq. (7) one can see that the upper bound of achievable precision is given by $1/\sqrt{t_{\mathrm{tot}}I^{Q}/t}$. This clearly suggests that the rescaled QFI $I^{Q}/t$ is the suitable figure of merit. For instance, in the case of many-body ground state sensing, one may consider an adiabatic approach for generating the complex ground state. This can be achieved by slowly evolving the system initialized in the ground state of a simpler Hamiltonian which is transformed into the desired Hamiltonian over the time interval $t$. If the transformation is slow enough, namely $t$ is large, the quantum state remains in the ground state of the instant Hamiltonian throughout the evolution. This guarantees that at the end of evolution when the Hamiltonian takes the final form and, hence, the system is in the desired ground state. When the system is adiabatically driven towards a second-order quantum phase transition, the required time $t$ scales as $t{\sim}N^{z}$ Rams et al. (2018), where $z$ is the dynamical critical exponent showing how the energy gap $\Delta E$ vanishes at the critical point by increasing the system size $\Delta E_{\rm min}\sim 1/N^{z}$. Incorporating this into the scaling of the QFI across second-order quantum phase transitions in Eq. (39) leads to
@@
This clearly shows that the scaling diminishes as time is included in the resource analysis. Note that the issue of time as a resource can arise in interferometric setups as well, for example in the adiabatic sweeping through quantum phase transitions for entanglement generation and detection Lee (2006); Huang et al. (2018); Zhuang et al. (2020).
@@
Similar results can be found for different quantum sensing strategies where the required time depends on the system size. For instance, in Ref. Montenegro et al. (2023) quantum-enhanced sensitivity, i.e. $\beta{>}1$, is achieved for the steady state at the boundary time crystal transition when the figure of merit is the QFI $I_{Q}$. However, if one changes their figure of merit to normalized QFI, i.e. $I_{Q}/t$, with $t$ being the time that the system reaches its steady state, the scaling with respect to system size is reduced to standard quantum limit. On the other hand, in the case of Stark systems, presented in section V.2, it has been shown that regardless of the number of particles and the range of the interaction, one always has $\beta-z{>}1$ ensuring that the quantum-enhanced sensitivity is always achievable for the Stark probes He et al. (2023); Yousefjani et al. (2023a).
Valuable insights on the normalized QFI can be obtained from the fundamental bounds on QFI. When the probe is the ground state of the Hamiltonian in the form $\hat{H}_{\theta}{=}\hat{H}_{0}(\theta){+}\hat{H}_{1}$, with a control term $\hat{H}_{1}$ and the parameter dependent term $\hat{H}_{0}(\theta)$, the upper bound of the QFI is given by Abiuso et al. (2025),
$I_{Q}(\theta)\leq\frac{||\partial_{\theta}\hat{H}_{0}(\theta)||^{2}}{\Delta^{2}},$ (63)
where the operator seminorm ||.||||.||| | . | | is the difference between the maximum and minimum eigenvalues of the operator, and $\Delta$ is the energy gap between the ground state and the first excited state. On the other hand, when the probe state is prepared by evolving an appropriate initial state with a Hamiltonian with a time-dependent control term in the form $\hat{H}_{\theta,t}{=}\hat{H}_{0}(\theta){+}\hat{H}_{1}(t)$, the QFI is upper bounded by the generalized Heisenberg limit Boixo et al. (2007); Puig et al. (2024)
$I_{Q}(\theta,t)\leq t^{2}||\partial_{\theta}\hat{H}_{0}(\theta)||^{2}.$ (64)
According to the adiabatic theorem, the time needed for adiabatic preparation of the ground state probe is inversely proportional to the square of the critical energy gap. However, with certain case specific shortcuts to adiabaticity, it can be made inversely proportional to the critical gap, namely $\sim 1/\Delta_{c}$ Roland and Cerf (2002). Therefore, the rescaled critical QFI in this case is bounded as $I_{Q}/t\leq\frac{||\partial_{\theta}\hat{H}_{0}(\theta)||^{2}}{\Delta_{c}}$. This indicates that the rescaled QFI can benefit from the scaling of both the energy gap as well as the seminorm factor. The fundamental bounds of the QFI in the context of different probes is provided in Table 3, which is adapted from Ref. Abiuso et al. (2025).
Encoding Maximum $I_{Q}(\theta)$
Dynamical $\rho_{\theta}=e^{-i\hat{H}_{\theta}t/\hbar}\rho e^{i\hat{H}_{\theta}t/\hbar}$ $\dfrac{t^{2}}{\hbar^{2}}\|\partial_{\theta}\hat{H}_{\theta}\|^{2}$
Thermal $\rho_{\theta}=\dfrac{e^{-\beta\hat{H}_{\theta}}}{\operatorname{Tr}{e^{-\beta\hat{H}_{\theta}}}}$ $\dfrac{\beta^{2}}{4}\|\partial_{\theta}\hat{H}_{\theta}\|^{2}$
Ground state of $\hat{H}_{\theta}$ with spectral gap $\Delta$    $\dfrac{1}{\Delta^{2}}\|\partial_{\theta}\hat{H}_{\theta}\|^{2}$
Table 3: QFI upper bounds for probes prepared dynamically, at thermal equilibrium (with inverse temperature $\beta$), and as ground states. Table adapted from Ref. Abiuso et al. (2025).
@@
The time needed for accomplishing the sensing task, is directly related to the strategy that one exploits. In Ref. Gietka et al. (2021), the authors show that by adiabatic evolution, one cannot reach Heisenberg scaling for quantum sensing. However, the proof relies on a specific choice of the adiabatic evolution. In their formalism the Hamiltonian is written as $\hat{H}{=}\theta\hat{H}_{0}{+}\hat{H}_{1}(t)$, where $\hat{H}_{0}$ and $\hat{H}_{1}$ do not depend on the target parameter $\theta$. In this scenario, $\hat{H}_{0}$ is always on and the time dependent Hamiltonian $\hat{H}(t)$ varies slowly. The resulted dynamics encodes the information about $\theta$ into the state of the system whose initial state does not depend on $\theta$. In this case, as proved in Ref. Gietka et al. (2021), Heisenberg scaling cannot be reached. Nonetheless, one may consider to prepare the probe in the ground state of $\hat{H}_{1}$ and as $\hat{H}_{1}$ is turned off, the probe is put under the action of a general $\hat{H}_{0}(\theta)$ adiabatically to prepare the probe in the ground state of a general $\hat{H}_{0}(\theta)$ instead of $\theta\hat{H}_{0}$. In this situation, obtaining Heisenberg scaling is possible.
@@ VIII Non-Equilibrium Quantum Sensing: Dynamical Quantum Phase Transitions
Figure 11: DQPT quantum sensor. (left) A Gaussian longitudinal magnetic field of the form $B_{x}(t)=B\exp\{-(t-\tau)^{2}/(2\sigma^{2})\}$ improves the Fisher information for measuring the coupling at $\tau=3t_{c_{1}}$. On the other hand for other $\tau$ and $B_{x}=0$ does not provide improvement in the Fisher information. The other parameters are $B=200$ G, $\sigma=\tau/2$, and $B_{z}=50G$ is fixed for both cases. Figure is adapted from González et al. (2022). (right) The scaling exponent of quantum Fisher information $I^{t\to\infty}_{Q}\sim N^{b}$ corresponding to $\Omega$ at the critical points $\Omega=\Omega_{c}$ of non-equilibrium phase transition. Abscissa represents the initial state for various choices of $\theta$ and $\phi=0$ of the spin-coherent state, $\omega/\chi=10^{-4}$, and $N\in[100,200]$. The lines are obtained numerically and dots are obtained using long-time secular contribution Pang and Brun (2014). Figure is adapted from Guan and Lewis-Swan (2021)
The footprints of the ground state quantum criticality in equilibrium can also be observed in non-equilibrium dynamics. For a closed system described by a time-independent Hamiltonian $\hat{H}$, its time evolution which is described by Schrödinger equation as
$|\Psi(t)\rangle=e^{-i\hat{H}t}|\Psi(0)\rangle,$ (65)
where $|\Psi(0)\rangle$ is the initial state. There are two different notions of dynamical quantum phase transitions, which we will discuss separately in the following subsections.
@@ VIII.1 Dynamical quantum phase transitions quantified by Loschmidt echo
Considering the projection of the state with the initial state that changed in time, one can define the Loschmidt echo, ${\cal L}(t)=|\langle\Psi(0)|\Psi(t)\rangle|^{2}$, interpreted as the probability of return. One can define the rate of return probability ${\cal G}$, as
${\cal G}(t)=-\lim_{N\to\infty}\frac{1}{N}\log{\cal}\mathcal{L}(t).$ (66)
It is discovered that if $|\Psi(0)\rangle$ and $|\Psi(t)\rangle$ belong to two different equilibrium phases, then ${\cal G}(t)$ can be non-analytic at a certain time $t{=}t_{c}$. The occurrence of non-analytic behavior has been rigorously established as the dynamical quantum phase transition (DQPT) Heyl (2018). On the other hand, if $|\Psi(0)\rangle$ and $|\Psi(t)\rangle$ belong to the same equilibrium phase, then ${\cal G}(t)$ is a regular function of time. There are a few peculiarities to these cases Jafari and Johannesson (2017) as well as their extension in non-Hermitian systems Zhou et al. (2018a); Jing et al. (2024). Such dynamical phase transitions have been experimentally realized in large quantum simulators (Zhang et al., 2017). Some recent efforts have been made to explore its relevance as a useful resource for various quantum technologies, including quantum-enhanced sensing González et al. (2022), where the authors have proposed a negatively charged Nitrogen-Vacancy (NV-) center in nearby Carbon-13 nuclear spins as a platform for studying dynamical phase transitions. First, they have shown that the system undergoes a dynamical phase transition by a suitable choice of a Gaussian time-varying magnetic field. Then it is shown that the Fisher information, for estimating the coupling strength between the two carbon-13 nuclear spins, improves for $\tau=3t_{c_{1}}$ as compared to the other choices of $\tau$ at regular times. This is shown in the left panel of Fig. 11.
@@ VIII.2 Dynamical quantum phase transitions quantified by a time-averaged order parameter
A second category of dynamical phase transitions in many-body systems is detected by the time-averaged order parameter ${\cal O}(t)$. The time average of the order parameter in the evolved state $|\Psi(t)\rangle$ is given by
@@
The order parameter distinguishes an ordered phase when $\overline{\langle{\cal O}(t)\rangle}\neq 0$ from a disordered phase when $\overline{\langle{\cal O}(t)\rangle}=0$. It has been shown that Fisher information can be a tool for identifying such steady state phase transitions Wu and Xu (2016); Prokopenko et al. (2011); Huang et al. (2016); Paz et al. (2024); Guan and Lewis-Swan (2021). Moreover, using a suitable collective spin model, steady-state phase transitions have been identified for quantum-enhanced sensing of single-parameter estimation Guan and Lewis-Swan (2021). In this paper, the authors have studied a fully connected spin model under sudden quenching. They focused on the LMG model which is given by
[…]
where $\hat{S}^{x,y,z}{=}\sum_{i=1}^{N}\hat{\sigma}^{x,y,z}_{i}/2$ are defined as collective spin operators and $\hat{\sigma}^{x,y,z}_{i}$ are the Pauli matrices for the $i$th spin-1/2 particle. The criticality of the ground state of the model occurs at $\Omega_{c}/\chi{=}0.5$ and $\omega_{c}/\chi{=}0$. The Fisher information in long-time state behaves as an order parameter and shows a peak at the critical point. Moreover, the scaling of the long-time QFI at the critical point as a function of system size $N$ reaches a sub-Heisenberg limit for the transverse field, $I^{t\to\infty}_{Q,\Omega}\sim N^{1.5}$ and $I^{t\to\infty}_{Q,\omega}\sim N^{1.75}$ for the longitudinal field. For various choices of initial state, the scaling exponent $b$, obtained by fitting the numerical data in the function $I_{Q}^{t\to\infty}\sim N^{b}$, is shown in the right panel of Fig. 11.
@@ IX Non-Equilibrium Quantum Sensing: Discrete Time Crystal Phase Transitions
Apart from dynamical phase transitions, discussed in Sec. VIII, there are other categories of phase transitions which occur through non-equilibrium dynamics of many-body systems. An important class of such systems are discrete-time crystals (DTCs) which are defined in an isolated non-equilibrium quantum many-body system that undergoes periodic drives. Breaking the transnational symmetry of time results in the emergence of a new phase of matter known as the DTC phase with persistent indefinite oscillations Sacha (2015); Khemani et al. (2016); Else et al. (2016); Yoshinaga et al. (2022). This concept has also been studied in the non-Hermitian systems Yousefjani et al. (2025c). Recently, long-range spatial and time-ordered dynamics in the DTC phase have been harnessed to measure an AC field Iemini et al. (2024b); Gribben et al. (2025); Yousefjani et al. (2025d). The optimal performance for long periods along with intrinsic robustness to the imperfections in the sensing protocol based on DTC phase allows one to hit the standard quantum limit in estimating the AC field. However, obtaining a true sensing enhancement based on a resource such as system size requires a wised mechanism for establishing DTC. In Ref. Yousefjani et al. (2025b) the proposed mechanism results in a stable DTC with period-doubling oscillations that persist indefinitely even in finite-size systems. The time-dependent Hamiltonian that governs the dynamic in the considered spin-$1/2$ chain is
$\hat{H}(t)=J\hat{H}_{I}+\sum_{n}\delta(t-nT)\hat{H}_{P},$ (69)
$\hat{H}_{I}=\sum_{j=1}^{L-1}j\hat{\sigma}^{z}_{j}\hat{\sigma}^{z}_{j+1},\quad\hat{H}_{P}=\Phi\sum_{j=1}^{L}\hat{\sigma}^{x}_{j}.$ (70)
@@
Here $J$ is the spin exchange coupling, and $\hat{\sigma}^{x,y,z}_{j}$ are the Pauli operators. For one period of evolution, the Floquet unitary operator reads
$\hat{U}_{F}(\Omega,\varepsilon)=e^{-i\hat{H}_{P}}e^{-i\Omega\hat{H}_{I}},$ (71)
@@
here $\Theta{=}JT$, and $\Phi$ is tuned to be $\Phi{=}(1{-}\varepsilon)\frac{\pi}{2}$, with $\varepsilon$ as deviation from a $\pi/2$ $x$-rotation. While setting $\Theta{=}\pi/2$ results in a stable period doubling DTC that is robust against arbitrary imperfection $\varepsilon$ in the rotating pulse, the system goes through a sharp second-order phase transition as the spin exchange coupling, namely $\Theta$ varies from $\frac{\pi}{2}$, denoting this variation as $\theta{=}|\Theta-\frac{\pi}{2}|$. Relying on this transition, a DTC quantum sensor has been devised to sense $\theta$ with benefits from multiple features. As it is clear from Fig. 12(a), in the DTC phase the QFI provides a plateau whose value depends on period cycles $n$, and in the non-DTC region it shows nontrivial and fast oscillations. By approaching the transition point, denoted by $\theta_{\max}$ (dashed line), the QFI shows a clear peak at all stroboscopic times. The extreme sensitivity to the exchange coupling across the whole DTC phase ($\theta{<}\theta_{\max}$) as well as at the transition point ($\theta_{\max}$) providing quantum-enhanced sensitivity as $I_{Q}{\propto}L^{\beta}$ with $\beta{>}3$ (see Fig.12 (b)). The obtained quantum enhancement is independent of the initial state and can be boosted by increasing imperfection in the pulse to a certain value.
@@
Figure 12: DTC quantum sensor; (a) The QFI versus $\theta$ in stroboscopic times in a system of size $L{=}30$. The onset of the phase transition is determined by $\theta{=}\theta_{\max}$ (dashed line), the point where QFI peaks in different period cycles $n$’s. (b) The values of the QFI after $n=2$ in DTC phase (for $\theta{=}10^{-4}\pi/2$) and at transition points ($\theta{=}\theta_{\max}$) versus $L$. The numerical simulation (NS) is well-mapped by a function as $I_{Q}{\propto}L^{\beta}$ (solid lines) with $\beta{>}3$. Results are obtained for $\varepsilon{=}0.01$. Figure is adapted from Yousefjani et al. (2025b).